# Deep Sparse Latent Feature Models for Knowledge Graph Completion

### v1

@@ A.4 Reparameterization
where $z_{hr,k}$ and $\pi_{hr,k}$ are defined in Equation 12 and 15, $\text{logit}(\cdot)$ is the inverse-sigmoid function and $\lambda$ is the relaxation temperature. The reparameterization of $z_{t,k}$ is achieved similarly.
@@
Table 7: The parameter count, training epochs, and GPU hours required by SimKGC (Wang et al., 2022) and DSLFM-KGC.
@@
Appendix B Experiment
B.1 Experiment settings
Datasets. The statistics of each dataset are shown in Table 5.
Baselines. The baseline methods we choose can be categorized into three classes:
  • For rule-based methods, we incorporate NeuralLP (Yang et al., 2017), DRUM (Sadeghian et al., 2019) and LERP (Han et al., 2023).
  • In the category of embedding-based methods, we choose TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), R-GCN (Schlichtkrull et al., 2018), ConvE (Dettmers et al., 2018), RotatE (Sun et al., 2019), TuckER (Balažević et al., 2019), HittER (Chen et al., 2021), N-Former (Liu et al., 2022) and KRACL Tan et al. (2023).
  • Text-based methods considered include KG-BERT (Yao et al., 2019), MTL-KGC (Kim et al., 2020), StAR (Wang et al., 2021a), KG-S2S (Chen et al., 2022), DKPL (Xie et al., 2016), KEPLER (Wang et al., 2021b), BLP Daza et al. (2021), SimKGC (Wang et al., 2022) and GHN (Qiao et al., 2023).
$x_{t}$ $=\left[\text{CLS},\mathcal{M}(t),\text{SEP}\right]$ (35)
where CLS and SEP are special tokens introduced by Devlin et al. (2019), and $\mathcal{M}(h),\mathcal{M}(r)$, and $\mathcal{M}(t)$ represent the tokenized textual descriptions of the head, relation and tail, respectively. Following tokenization, $\mathbf{x}_{hr}$ and $\mathbf{x}_{t}$ are processed through BERT encoders, as specified in Equation 14.
@@ B.1 Experiment settings
(a) WN18RR
(b) FB15k-237
Table 9: Performance of DSLFM-KGC on the WN18RR and FB15k-237 datasets w/ different $\beta$ values.
@@
B.2 Additional ablation results
@@
Appendix C Related work
@@
KGC with large language models (LLMs). Recent advancements in text-based KGC leverage the extensive pre-trained knowledge and contextual understanding of LLMs to bridge the gap between structured and unstructured knowledge. Techniques in this domain often employ diverse prompt designs to enable LLMs to perform direct reasoning for KGC (Yao et al., 2023; Wei et al., 2024) or to refine textual information in datasets, enhancing their accuracy and richness (Li et al., 2023a; Yang et al., 2024). However, while these methods are training-free and inherently interpretable, they face challenges such as hallucinations and reliance on few-shot demonstrations, which are difficult to implement in sparsely connected KGs like WN18RR. Alternatively, some approaches fine-tune LLMs on KGC tasks using strategies like prefix-tuning (Chen et al., 2023; Zhang et al., 2024b) or adapter-tuning. While these methods capitalize on the reasoning capabilities of LLMs, they often lack interpretability, struggle to generalize across datasets, and continue to face challenges in achieving strong performance. In contrast, our model excels on relatively sparse KGs with distinct clustering patterns, leveraging text not only to improve KGC interpretability but also to provide meaningful clustering information about the KG itself. Additionally, while LLMs provide external knowledge to enhance KGC, our approach focuses on directly extracting and utilizing the intrinsic information within KGs to strengthen representation learning. This makes our method particularly effective in scenarios where LLMs cannot reliably provide external knowledge, such as in domain-specific datasets.
@@
Our work also relates closely to Variational AutoEncoders (VAEs) (Kingma & Welling, 2013), a foundational class of generative models that employs an encoder to map input data to a latent space, typically assuming a Gaussian prior, and a decoder to reconstruct the data from this latent representation. To facilitate gradient-based optimization during training, the reparameterization trick is used, re-expressing the sampling of latent variables as deterministic functions of noise variables, thereby enabling backpropagation through stochastic nodes. While this trick is straightforward for "location-scale" distributions like the Gaussian, extending it to other distributions such as Bernoulli (Jang et al., 2016; Maddison et al., 2016) and Beta distributions (Nalisnick & Smyth, 2016) requires more sophisticated techniques. Reparameterization for these distributions often involves implicit differentiation methods to compute gradients when explicit reparameterization is infeasible (Figurnov et al., 2018). A persistent challenge in training VAEs is posterior collapse, where the encoder’s output becomes similar to the prior, causing the model to ignore the latent variables (Bowman et al., 2015). This issue undermines the VAE’s ability to learn meaningful representations. Various strategies have been proposed to mitigate posterior collapse, including modifying the objective function with $\beta$VAE to balance reconstruction and regularization terms (Higgins et al., 2017), employing annealing schedules for the KL divergence term (Bowman et al., 2015), and designing more expressive posterior distributions to better capture the underlying data structure (Rezende & Mohamed, 2015).

### v2

@@ B.4 Reparameterization
where $z_{hr,k}$ and $\pi_{hr,k}$ are defined in Equation 13 and 16, $\text{logit}(\cdot)$ is the inverse-sigmoid function and $\lambda$ is the relaxation temperature. The reparameterization of $z_{t,k}$ is achieved similarly.
Appendix C Experimental Details
@@
(a) WN18RR
(b) FB15k-237
Datasets. The statistics of each dataset are shown in Table 3. The UMLS encompasses a wide array of biomedical concepts, including names for drugs and diseases, alongside the relationships that exist between them, such as diagnosis and treatment. Originally introduced by [5], the WN18 and FB15k datasets were later refined to WN18RR and FB15k-237 following studies [51, 13] that revealed test leakage issues. Textual data comes from KG-BERT [62]. Wikidata5M-Ind [56] is a large-scale KG merging Wikidata and Wikipedia, with textual descriptions for each entity.
Table 5: The parameter count, training epochs, and GPU hours required by SimKGC [55] and DSLFM-KGC.
@@
Baselines. The baseline methods we choose can be categorized into four classes:
  • In the category of embedding-based methods, we choose TransE [5], DistMult [59], ConvE [13], RotatE [48], TuckER [3], HittER [9] and KRACL [49].
  • For GNN-based methods, we consider R-GCN [45], CompGCN [52] and KGGAT [41].
  • Text-based methods considered include KG-BERT [62], MTL-KGC [26], StAR [54], KG-S2S [7], DKPL [58], KEPLER [56], BLP [12], SimKGC [55] and GHN [43].
  • For the LLM-based methods, we include CP-KGC [60] and KICGPT [57].
$x_{t}$ $=\left[\text{CLS},\mathcal{M}(t),\text{SEP}\right]$ (35)
where CLS and SEP are special tokens introduced by [14], and $\mathcal{M}(h),\mathcal{M}(r)$, and $\mathcal{M}(t)$ represent the tokenized textual descriptions of the head, relation and tail, respectively. Following tokenization, $\mathbf{x}_{hr}$ and $\mathbf{x}_{t}$ are processed through BERT encoders, as specified in Equation 15.
Figure 7: Testing data distribution of the WN18RR and FB15k-237 dataset w.r.t. neighborhoods at varying geodesic distances, where ‘5+’ indicates a distance greater than 5.
Figure 8: Comparison of DSLFM-KGC and SimKGC performance on the FB15k-237 dataset w.r.t. neighborhoods at varying geodesic distances, where ‘5+’ denotes a distance greater than 5. If no path exists between two entities, the distance is considered infinite.
Appendix D Additional results
Figure 9: Performance of DSLFM-KGC on the WN18RR and FB15k-237 datasets w/ different $\beta$ values.
@@ Appendix D Additional results
D.1 Additional ablation results
@@
D.2 Additional analysis results
Figure 10: Modularity score of the WN18RR and FB15k-237 datasets.
Dataset Modularity score
WN18RR 0.574
FB15k-237 0.074
Modularity of the KGs. To assess the significance of clustering within the WN18RR and FB15k-237 datasets, we quantified the modularity of each. Specifically, we applied the label propagation algorithm for community detection and subsequently calculated the modularity score of the resulting partitions using the following equation:
$Q=\frac{1}{2m}\sum_{i,j}\left(A_{ij}-\gamma\frac{k_{i}k_{j}}{2m}\right)\delta(c_{i},c_{j})$ (36)
where $m$ is the number of edges, $A$ is the adjacency matrix, $\gamma$ is the resolution parameter, $k_{i}$ and $k_{j}$ are the degrees of nodes $i$ and $j$, and $\delta(c_{i},c_{j})$ is 1 if nodes $i$ and $j$ belong to the same community, and 0 otherwise. A modularity score closer to 1 indicates a more distinct clustering structure.
Geodesic analysis on the FB15k-237 dataset. Figure 8 presents the performance comparison between DSLFM-KGC and SimKGC w.r.t. geodesic distance on the FB15k-237 dataset. To determine the triple distribution, we first construct the graph using training data and subsequently compute the shortest path distance between the head and tail entity for each triple.
Visualization of the latent structure. For demonstration purposes, we use stick-breaking prior settings of $\alpha_{\text{qry}}=100$, $\alpha_{\text{ans}}=50$, and a truncation level of $K=64$ to generate the sparse latent feature matrix $F_{\text{ans}}$. In Figure 11, each matrix shows how entities are grouped into communities, where larger absolute values suggest stronger confidence in whether a node belongs to a specific community.
(a) WN18RR
(b) FB15k-237
Cluster : County
County Wexford: County Wexford is a county in Ireland…
Cluster: Music
PJ Harvey: Polly Jean Harvey MBE is an English musician…
Italo disco: Italo disco is a genre of music which originated in Italy…
Talent manager-GB: A talent manager, also known as band manager…
Appendix E Related work
@@
KGC with large language models (LLMs). Recent advancements in text-based KGC leverage the extensive pre-trained knowledge and contextual understanding of LLMs to bridge the gap between structured and unstructured knowledge. Techniques in this domain often employ diverse prompt designs to enable LLMs to perform direct reasoning for KGC [63, 57] or to refine textual information in datasets, enhancing their accuracy and richness [33, 60]. However, while these methods are training-free and inherently interpretable, they face challenges such as hallucinations and reliance on few-shot demonstrations, which are difficult to implement in sparsely connected KGs like WN18RR. Alternatively, some approaches fine-tune LLMs on KGC tasks using strategies like prefix-tuning [8, 65] or adapter-tuning. While these methods capitalize on the reasoning capabilities of LLMs, they often lack interpretability, struggle to generalize across datasets, and continue to face challenges in achieving strong performance. In contrast, our model excels on relatively sparse KGs with distinct clustering patterns, leveraging text not only to improve KGC interpretability but also to provide meaningful clustering information about the KG itself. Additionally, while LLMs provide external knowledge to enhance KGC, our approach focuses on directly extracting and utilizing the intrinsic information within KGs to strengthen representation learning. This makes our method particularly effective in scenarios where LLMs cannot reliably provide external knowledge, such as in domain-specific datasets.
@@
Our work also relates closely to Variational AutoEncoders (VAEs) [27], a foundational class of generative models that employs an encoder to map input data to a latent space, typically assuming a Gaussian prior, and a decoder to reconstruct the data from this latent representation. To facilitate gradient-based optimization during training, the reparameterization trick is used, re-expressing the sampling of latent variables as deterministic functions of noise variables, thereby enabling backpropagation through stochastic nodes. While this trick is straightforward for "location-scale" distributions like the Gaussian, extending it to other distributions such as Bernoulli [23, 36] and Beta distributions [40] requires more sophisticated techniques. Reparameterization for these distributions often involves implicit differentiation methods to compute gradients when explicit reparameterization is infeasible [16]. A persistent challenge in training VAEs is posterior collapse, where the encoder’s output becomes similar to the prior, causing the model to ignore the latent variables [6]. This issue undermines the VAE’s ability to learn meaningful representations. Various strategies have been proposed to mitigate posterior collapse, including modifying the objective function with $\beta$VAE to balance reconstruction and regularization terms [20], employing annealing schedules for the KL divergence term [6], and designing more expressive posterior distributions to better capture the underlying data structure [44].